# Y-AR: A Mixed Reality CAD Tool for 3D Wire Bending

### v1

@@ 5. Design of Y-AR
Y-AR follows a verb-noun interaction model. The Y-AR interface consists of a top-level menu with different actions: creating connectors, mid-air drawing, and creating marble tracks. Next to that, it comes with basic editing tools such as linking structures, deleting, and modifying or moving components and a generic export function. As illustrated by Figure 2, Y-AR allows users to create models that are hard to bend by hand and challenging to produce with CAM tools.
@@ 5.1. Create connector tool
The create connector tool lets designers create wire-bending structures to connect to existing objects or to features in the context of the user. The workflow for this tool is as follows:
Users select a connector type and press the gesture button to enable gesture detection (this is not ”always on” to avoid undesired detection of gestures). They simply grab the object the way they want to connect, Y-AR detects the grabbing gesture and extracts dimensions for the required connector. Dimensions include the distance between the index finger and thumb to control the size of the generated connector, the hand’s orientation, and the palm’s position to control the placement of the generated connector, find specific details in section 7.1.
The generated connector shows up in the XR model as a preview of what the wire structure will look like. Upon placement, the modify tool is triggered so users can immediately move, rotate and scale the connector as needed.
As mentioned before, Y-AR requires users to choose the connector type. This allows it to identify the right parameters for each type of connector. Below we specify how these connector types relate to their respective gestures.
The table edge connector measures the table edge’s thickness using the vertical distance between the thumb tip and palm (see Figure 5 b).
The cylindrical connector determines the cylinder diameter based on the distance between the thumb and index finger during a grabbing gesture (see Figure 5 c).
And the pegboard connector sets the distance between connector legs using the spread of the index and middle fingers (see Figure 5 a).
@@ 5.2. Mid-air drawing tool
The mid-air drawing tool allows users to freely draw lines in mid-air, generating paths with more flexibility. As shown in Figure 6, users can draw in various contexts, such as (a) mid-air, (b) on the human body, (c) on the surfaces of existing objects like lampshades for decorative purposes, or (d) create mesh structures based on dress forms. These capabilities enable the creation of unique wire structures, such as wire-wrapped jewelry, which benefit from lightweight connections, high flexibility, and aesthetic customization.
In our implementation of Y-AR, users use the Logitech MX Ink, pen-based tool for Mixed Reality (MR) with high-accuracy drawing functions, to start drawing a line the user press the button in the MX Ink and draws in mid-air, each line end upon releasing the button. Alternatively, users perform a pinch gesture to start drawing a line until the pinch gesture is ended without the need of the MX Ink. Users freely draw lines to model their desired structures in 3D.
@@ 5.3. Marble track tool
To create more deliberate functional structures, we let users design structures using what we call Shape Modules. To exemplify, we implemented a Marble Track tool demonstrated in Figure 7. Users create freehand structures consisting of pre-defined modules. This tool is developed based on the Mid-Air Drawing Tool, and its basic workflow is similar.
@@
When users sketch a line for a marble track, the system automatically generates not just a single wire, but four parallel lines to form a robust track. It occasionally injects support towers consisting of vertical truss structures as well as small 3D printed components to hold the wires together. Additionally, based on the track’s length, Y-AR places additional support structures at midpoints along the track to prevent sagging and maintain structural integrity under the weight of the marbles. Our current marble track modules are optimized for standard 1.6 cm in diameter marbles, it would be relatively easy to extend to other sizes and types.
5.4. Other editing tools
@@
Additionally we provide a set of generic editing tools to modify designs. These tools apply to all structures and content in the scene. They consist of a delete tool to delete segments or lines; a modify tool, to move scale or rotate segments or lines, and a link tool to link individual line segments or connectors to one another. Finally, there is one export function to convert the designed wires to machine instructions (csv) and assembly instructions (AR preview).
@@ 5.5. Gesture training
Traditionally, gesture-based interfaces are not the most discoverable (Zhou et al., 2016). However, to create connectors, Y-AR leverages gesture interaction by having users demonstrate how the structure should mount objects. In doing so, we build on work by Pei et al. (2022), which lets users’ hands become virtual objects by imitating the objects themselves. To aid users in adopting these gestures, every tool in Y-AR includes an interactive on-boarding mode, as shown in Figure 8.
@@
Y-AR is a Unity application, it implements 3D and UI interactions via Unity input events and includes customizable 3D UI elements for on-boarding, spatial data collection, and Passthrough mode setup. The chosen tech stack, including Unity version 2022.3.8f1, XR Interaction Toolkit (version 2.5.1), XR Hands (version 1.4.1), XR Core Utilities (version 2.2.3), XRC Line Tool (version 0.1.0), and XR Core (version 0.1.4), was selected for its comprehensive support for mixed reality development and ease of integration with various XR features.
@@
Y-AR achieves hand gesture tracking by using the Unity XR Hands package. Gesture recognition translates hand movements into CAD system actions, with real-time feedback provided by the XRC Line Tool from the XRC core, which is used for creating, rendering, and manipulating lines in the project.
@@
Y-AR achieves the transformation of spatial data into a wire-bending path through a dynamic tracing system that accurately places and scales connectors. This system measures distances between specific hand joints to proportionally adjust the sizes of virtual objects, ensuring precise and adaptable designs. Y-AR achieves mid-air drawing by using the controller’s position as input, visualizing this with the XRC Line Tool. The system utilizes compatible controllers to track spatial movements, which are then translated into wire-bending paths and exported as drawing data.
@@
The bending paths designed in Y-AR are exported as CSV files, containing a list of coordinates along the path. This format is chosen for its simplicity and compatibility, allowing the files to be easily read and interpreted by wire bending software. This ensures that the designs can be accurately translated into physical forms by the wire bender.
7. Fabrication-Aware Design for Wire Bending
In this section, we systematically explore fabrication-aware design for wire bending through the design of different types of connectors, and the limitations of materials and fabrication machines.
7.1. Connector Design
We designed a series of connectors using wire bending to connect different types of objects. As shown in Figure 9, the target objects are defined as collective variances of geometric primitives (including: (a) a hanging rectangular shape, (b) a cylindrical shape supported on the bottom, (c) a hanging cylindrical shape, (d) a horizontally oriented rectangle, and (e) spherical shapes).
Based on these designs, we developed several design guidelines to guarantee successful wire bending with a numerical bending machine: (1) avoiding planar bends, (2) minimum bending angle to overcome plastic deformation, (3) constraining 3DoF, (4) collision detection.
7.2. Avoid planar bends
In our design tests, we often find that when performing multiple bends (e.g., three consecutive 90-degree bends), the wire may collide with the machine or the extruded wire itself, leading to wire bending failure. To address this issue, our strategy is to introduce rotations at the nodes, particularly when designing planar structures as shown in Figure 10 a. By incorporating rotations, as illustrated in Figure 10 b, we ensure that previously bent sections of the wire and subsequent bends are not on the same plane, thereby avoiding collisions and ensuring successful wire bending.
7.3. Minimum bending angle to overcome plastic deformation
To plastically deform wires (bend them permanently), it is crucial to overcome the elastic regime of the material. Small bends simply bend the wire and the wire bends back. To overcome this, we identify a minimum required bending angle and over-extend bends where needed.
In our design process, we tested bending angles from 5, 15, 25, 35, 45… up to 85 degrees (within a 90-degree range) and compared these with the actual bending results. This approach helped us effectively avoid the issue of ensuring that small bending angles in the design would still result in a noticeable bend.
As shown in the Figure 11, when the machine was set to bend at 15 degrees, the actual bending angle was significantly reduced. This finding indicates that in our designs, we should avoid specifying angles smaller than or around 15 degrees.
7.4. Constraining 3DOF
In our connector design, we primarily focus on constraining the three degrees of freedom (minus rotation) of objects. This means that to fix an object in a spatial position, all planar degrees of freedom need to be restricted. We use three main strategies to achieve this:
Gravity-Based Holders. These connectors rely on gravity to hold objects in place. As shown in Figure 9 c, cylindrical objects, such as water bottles, hang on the holder using their weight. The drawback of this method is that if the holder’s orientation changes, the fixed object may no longer be secured.
Clamp-Type Holders. These connectors use force in one direction to secure the object, while the object’s inherent limitations prevent movement in another direction. Examples include Figures 9 a and d. The advantage of this method is that, ideally, if the force is strong enough, the object remains fixed even if the connector’s orientation changes. However, as seen in Figure 12 a, if the force (both positive and negative Y direction) is insufficient, the object might move along the X direction perpendicular to the applied force.
Combination Holders. These connectors combine the principles of both gravity-based and clamp-type holders. An example is Figure 9 b (cafe cup holder) and Figure 9 e. The cafe cup holder clamps the coffee cup to prevent it from spilling, applying force in both positive and negative Y-directions (As shown in 12 b). Gravity effectively restricts movement in the vertical direction. This combined approach ensures stability and minimizes the risk of the object becoming unsecured due to changes in orientation.
7.5. Collision detection
Users lack knowledge about the machine’s limitations to bend wires. Issues can arise when lines are too close to each other or when the angle between them is too small, potentially causing machine collisions during bending.
To address this, we implemented basic collision detection. Our algorithm sets a minimum threshold angle that the wire bender can fabricate. Since all lines are drawn by connecting multiple points, the algorithm separates all points and combines every pair of points into a line segment. By calculating the angles between two such segments lines, the algorithm determines whether their angle exceeds the threshold. If it is, those segments are flagged as conflicts.
@@ 8. USER STUDY
To validate our implementation of Y-AR we conduct a basic usability study, demonstrating that non-technical users can indeed reliably design and fabricate mounting structures with Y-AR.
8.1. Participants
We recruited 12 participants from our university (number of female students = 7, age mean = 24.33, SD = 1.83). Most participants (10) reported having beginner proficiency with CAD tools, and only a few (3) indicated they engaged in wire bending activities before.
8.2. Hardware
We used the Meta Quest 3 as the study apparatus. We fabricated the designed structured during the study with the C64 CNC Wire Bender from PensaLabs. As input controller, we used the MX INK from Logitech.
8.3. Task
Each participant performed two tasks with Y-AR. Task 1 involves designingsm to connect a bottle to a table edge (see Figure 14). Task 2 is an open-ended exploration of Y-AR.
8.4. Metrics
To assess the usability, we selected specific questions from standardized tests: the System Usability Scale (SUS) and the User Experience Questionnaire (UEQ). Additionally, the mechanical performance of the resulting structures was evaluated by mounting a bottle to determine whether it could be securely supported.
8.5. Procedure
This study lasted 60 minutes in a controlled lab setting. Participants performed two tasks and were given 5 minutes to familiarize themselves with the software through a training video. In the first 15-minute task, they conducted task 1. If participants did not complete their modeling within the allocated time, their modeling was marked as incomplete. If they finished early, they could inform the researcher and proceed to the next stage. They fabricate the structure with help from the researchers and assemble their designs. They then proceeded with task 2 in another 15-minute block.
Finally, participants filled out our usability survey and provided feedback on their experiences in a semi-structured interview. All participants completed the study within 60 minutes.

### v2

@@ 6. Design of Y-AR
Y-AR adopts a verb-noun interaction model to make actions explicit and easy to discover, helping users understand what they can do (verbs) and what objects they can act upon (nouns). The Y-AR interface consists of a top-level menu with various actions, including creating connectors, mid-air drawing, and creating marble tracks. For basic editing, users can link structures, delete components, or move them. To transition to fabrication, it has an export function. In this section, we discuss the design rationale behind these tools.
@@ 6.1. Create Connector Tool
The Create Connector tool allows designers to create wire-bending structures that connect to existing objects or features within the user’s context. The workflow for this tool is as follows:
Users first press the gesture button to enable gesture detection (note that this feature is not ”always on” to prevent undesired detection of gestures). They grab the object the way they want to connect, Y-AR detects the gesture, and extracts dimensions to scale the connector. For example, dimensions may include the distance between the index finger and thumb to control the size of the generated connector, the hand’s orientation, and the palm’s position to control placement, depending on the type of gesture. The generated connector shows up as a preview of the wire structure. Upon placement, the modify tool is triggered, allowing users to immediately move, rotate, and scale the connector as needed.
Y-AR’s gestures intentionally mimic how a wire structure would physically interact with an object. For instance, a pinching gesture naturally corresponds to a clamping connector; a grab posture resembles wrapping a cylinder. This embodied mapping helps users intuitively specify both the intent and form of the connector during creation.
Below, we provide a more detailed description of how these connector types correspond to their respective gestures. (see Figure 8)
@@ 6.2. Mid-Air Drawing Tool
The mid-air drawing tool lets users generate paths with more flexibility. As shown in Figure 9, users can draw in various contexts, such as (a) mid-air, (b) on the human body, (c) on the surfaces of existing objects, like lampshades for decorative purposes, or (d) to create mesh structures based on dress forms. These capabilities enable the creation of unique wire structures, such as wire-wrapped jewelry, which benefit from lightweight connections, high flexibility, and aesthetic customization.
Y-AR, users employ the Logitech MX Ink, a pen-based tool for Mixed Reality (MR), to draw lines in mid-air.
By pressing the button on the MX Ink, users begin drawing; each line ends upon releasing the button. Alternatively, users perform a pinch gesture to start drawing a line, which continues until the gesture ends—eliminating the need for the MX Ink. Users can freely draw lines to model their desired structures directly in 3D space.
@@ 6.3. Marble Track Tool
To create more deliberate functional structures, we let users design structures using the create marble track tool, demonstrated in Figure 10. Users create freehand structures consisting of pre-defined modules. This tool is developed based on the Mid-Air Drawing Tool, and its basic workflow is similar.
@@
When users sketch a line for a marble track, Y-AR automatically generates not just a single wire, but four parallel lines to form a robust track. It injects support towers, consisting of vertical truss structures, at the starting point of each drawn wire segment to provide an anchor for stability. Additionally, 3D-printed connectors are placed along the wire’s length at every 5 cm to secure the wires together and ensure structural rigidity. Our current marble track modules are optimized for standard 1.6 cm diameter marbles, and the geometric parameters used in track generation (such as minimum bend radius and spacing) follow the same machine-specific simplification logic described in Section Simplification Parameters.
6.4. Editing tools
@@
Y-AR includes a set of simple editing tools. These tools can be applied to all structures and content in the scene. They consist of a delete tool to remove segments, a modify tool to move, scale, or rotate segments, and a link tool to link individual connectors to one another. Finally, there is an export function that converts the designed wires to machine instructions (CSV format) and assembly instructions (AR preview).
6.4.1. Simplification Parameters
Geometric processing steps in Y-AR are informed by the physical constraints of the bending machine and material. When simplifying user-drawn lines, we apply smoothing and point reduction based on a minimum bendable angle and a minimum spacing between bends, ensuring that generated geometries remain feasible for fabrication. Additionally, we expose both smoothing strength and minimum point reduction ratio as adjustable settings, allowing users to control the degree of simplification.
These parameters are computed from a simple geometric analysis of the bending machine’s head. Specifically, we calculate the minimum distance required between adjacent bends by measuring the die diameter (D1), pin diameter (D2), wire thickness (D3), and an optional clearance gap (G), as shown in Fig. 11a. The minimum bendable angle is derived from these measurements using the following relations:
Fig. 11b shows a photograph of the wire bender head used in our setup. These measurements ensure that simplification algorithms remain sensitive to the physical limitations of different wire benders and wire diameters. While currently configured for 1.6mm aluminum wire and a specific bending head, these constants can be easily adjusted to accommodate other materials and machines.
6.4.2. Collision detection
After simplification, Y-AR analyzes the resulting paths for potential fabrication issues. It sets a minimum threshold angle that the wire bender can fabricate. Since all lines are drawn (as shown in the Fig 12 a) by connecting multiple points, the algorithm separates all points and combines every pair of points into a line segment.
By calculating the angles between two such segment lines, the algorithm determines whether their angle exceeds the threshold. Conflicting segments are flagged, and Y-AR visualizes them by drawing short green highlight lines directly between the closest points on the conflicting segments using a distinctive material (as shown in the Fig 12 b). This provides an immediate visual indication of where the bending machine may encounter issues. It serves as an informative prompt, nudging users to make manual adjustments. While future work could enable automatic path adjustments or suggest alternatives, such functionality falls outside the scope of this paper.
6.4.3. Machine instructions and assembly instructions export
When users finish their design, Y-AR exports the wire structures as CSV files and continues to display a preview in AR to assist with assembly.
For connector-type structures, Y-AR generates separate CSV files for the connector ends and the wire segments between them. These parts are labeled directly on the virtual objects (Fig. 13 a), making it easy to match exported files with assembly steps.
For structures created using the Line Drawing Tool, the system exports one CSV file per line segment. Each line receives a number label in AR as soon as it is drawn (Fig. 13 b), so even complex designs with many lines are easy to trace.
After export, Y-AR keeps the numbered preview visible in AR, so users can clearly see where every part should go and assemble their design accurately.
@@ 6.5. Gesture training
Gesture-based interfaces are not the most discoverable (Zhou et al., 2016). Yet, we believe the intuitive nature of mimicking connector forms with hand gestures offers a clear benefit to end users. In doing so, we build on work by Pei et al. (2022), which lets users’ hands become virtual objects by imitating the objects themselves. To aid users in adopting these gestures, every tool in Y-AR includes an interactive on-boarding mode, as shown in Figure 14. This is only triggered upon the first usage of the tool, users can reset them so a different user on the same account can see on-boarding tutorials again.
@@ 7. Implementation
Y-AR is implemented in Unity3D (C#). It implements 3D and UI interactions via Unity input events and includes customizable 3D UI elements for on-boarding, spatial data collection, and Passthrough mode setup. The chosen tech stack, including Unity version 2022.3.8f1, XR Interaction Toolkit (version 2.5.1)888https://docs.unity3d.com/Packages/com.unity.xr.interaction.toolkit@2.5/manual/, XR Hands (version 1.4.1)999https://docs.unity3d.com/Packages/com.unity.xr.hands@1.4/manual/, XR Core Utilities (version 2.2.3)101010https://docs.unity3d.com/Packages/com.unity.xr.core-utils@2.2/manual/index.html, and XRC Line Tool (version 0.1.0)111111https://xrcollaboratory.github.io/edu.cornell.xrc.tools.line/, was selected for its comprehensive support for mixed reality development and ease of integration with various XR features.
@@
Y-AR uses the Unity XR Hands package for hand tracking. Gesture recognition translates hand movements into CAD system actions, with real-time feedback provided by the XRC Line Tool, which is used for creating, rendering, and manipulating lines in the project.
@@
Y-AR supports mid-air drawing and dynamic placement/scaling of virtual connectors by using XR Interaction Toolkit input events and XR Hands joint positions. We implemented additional logic for recording spatial data and real-time object scaling, which is visualized using the XRC Line Tool. Y-AR achieves mid-air drawing by using the controller’s position as input, visualizing this with the XRC Line Tool. The system utilizes compatible controllers to track spatial movements, which is then translated into wire-bending paths and exported as drawing data.
@@
The bending paths designed in Y-AR are exported as CSV files, which contain a list of coordinates along the path. This format is chosen for its simplicity and compatibility, allowing the files to be easily read and interpreted by wire bending software. This ensures that virtual designs are accurately translated into physical wire forms.
8. Technical Evaluation
We run two technical evaluations to validate the utility of our technique: the structural integrity of our basic connectors and the accuracy of bending angles. These findings inform the implementation of Y-AR. In our evaluations, we use 1.6mm aluminum wire, a commonly available material for wire bending. We use this fixed material to characterize the general behavior and efficacy of our technique, given that our contribution is not the calibration of wire-bending machines.
8.1. Pulling Force Resistance in Spring Wire Connectors
Y-AR leverages the inherent springiness of bent wire structures, which enables secure attachment, even when fabricated with small geometric inaccuracies. To evaluate this property, we conducted a mechanical test using a representative connector designed to hold a paper cup on the edge of a table.
We fabricated the connector using 1.6 mm aluminum wire and incrementally increased the load by filling the cup with water in 50 gram increments. This process simulates an increasing downward pulling force, and we define failure as the point at which the structure can no longer hold the cup or when water starts to spill.
In this test, water spills out when the load reaches 351 grams (350ml of water and the weight of the cup), as shown in Figure 15c. Even as water begins to spill, the structure continues to hold the cup securely. The structure displays only moderate elastic deformation (as shown in the Fig 15d ). The sequence of loading stages illustrates the progressive deformation leading up to this point, as shown in Figure 15b-c.
We conclude that the connector does fulfill its purpose until the cup is effectively full, to connect heavier objects, further reinforcement or thicker wire will be required. The software indicates an approximation of the weight connectors can hold. Future work could automatically reinforce when that value is exceeded.
8.2. Assessing Bending Accuracy for Small Angles
Wire bending requires overcoming both elastic and plastic deformation, as a result of this, simply bending to a certain angle may not suffice as the elastic deformation (or springiness) will make the material bounce back to some extend. Here, we characterize the bending behavior compared to our target angles.
We evaluate bending angles from 5∘, 15∘, 25∘, 35∘, 45∘, in 10∘ increments up to 85 degrees and compare to measured angles.We repeated this test five times and observed consistent results, Figure 16 shows one representative trial.
One critical result is that for small angles, a significant over-bend is required, a 15∘ of less bend will not achieve plastic deformation. In Y-AR this finding is used to avoid short bends with small angles, and if required by design they are accomplished by over-bending past the 15∘ mark. For larger angles some over-bending is still required, but to a lower extend.
@@ 9. User Study
We recruited participants to design using Y-AR. To assess the usability of Y-AR, we analyze their reactions and feedback.
Participants.
We recruited 12 participants (9female/3male/0 other; mean age = 21.4, SD = 4.0) from our university (see Table 1). We asked participants how frequently they used CAD tools and XR. P1, P3, P6, P7, P9, P11, and P12 reported prior exposure to wire-bending.
Apparatus.
Participants used a Meta Quest 3s headset, with the Logitech MX INK stylus as input controller. All fabricated wire structures were fabricated on the PensaLabs C64 CNC Wire Bender121212https://www.pensalabs.com/c64, with 1.6mm aluminum wire.
Training.
Participants received a brief on-boarding session, including headset setup, a built-in tutorial on hand gestures, and a short demonstration video showing key Y-AR features and a complete modeling example.
Task 1: Participants were tasked to create a functional object in response to a clearly defined real-world need. Specifically, they were asked to design a wire connector that would attach a standard 236 ml plastic bottle to the edge of a flat table (as shown in Fig. 17).
Before designing, participants were encouraged to think about how they might physically secure the bottle to the table edge using their own hands, and then to explore corresponding gestures in Y-AR to design their connector.
Task 2: An open ended task using Y-AR’s freehand design tool. Participants were encouraged to explore the surrounding environment—such as tables, chairs, shelves, and cables—and construct wire structures that interact with these elements. During the process, participants were encouraged to discuss their design decisions aloud and iteratively model their ideas using Y-AR.
Procedure
Each study session lasted approximately 60 minutes and was conducted in a controlled laboratory environment.
After training, participants proceeded to Task 1, with up to 20 minutes to design a functional wire connector. Upon completing the design, participants filled out a System Usability Scale (SUS) questionnaire reflecting on their experience.
They then proceeded to Task 2 for 15 minutes. At the end of the session, participants took part in a short semi-structured interview to discuss their design decisions, challenges, and overall impressions of Y-AR.
Metrics
We collected both subjective and objective metrics:
System Usability Scale (SUS): Participants rated the overall usability of Y-AR using the 10-item survey (Brooke, 1996).
Mechanical success: We evaluated whether the fabricated wire structures could securely perform their intended function (e.g., supporting a bottle on a table edge).
Semi-structured interviews: We analyzed qualitative feedback collected from post-task semi-structured interviews.